A soccer player starts from rest and accelerates forward, reaching a velocity of 8.00m/s in 2.50s. a) What was his average acceleration? b) what average force did he exert backward on the ground to achieve this acceleration? The players mass is 70.0kg, and air resistance is negligible
1
Expert's answer
2019-07-01T09:26:52-0400

a)


"a_{avg}=\frac{\Delta v}{\Delta t}=\frac{8}{2.5}=3.20\frac{m}{s^2}"

b)


"F=ma_{avg}=(70)(3.2)=224\ N"